# Automated Generation of Challenging Multiple-Choice Questions for Vision Language Model Evaluation

### v1

@@ Abstract
The rapid development of vision language models (VLMs) demands rigorous and reliable evaluation. However, current visual question answering (VQA) benchmarks often depend on open-ended questions, making accurate evaluation difficult due to the variability in natural language responses. To address this, we introduce AutoConverter, an agentic framework that automatically converts these open-ended questions into multiple-choice format, enabling objective evaluation while reducing the costly question creation process. Our experiments demonstrate that AutoConverter can generate correct and challenging multiple-choice questions, with VLMs demonstrating consistently similar or lower accuracy on these questions compared to human-created ones. Using AutoConverter, we construct VMCBench, a benchmark created by transforming 20 existing VQA datasets into a unified multiple-choice format, totaling 9,018 questions. We comprehensively evaluate 33 state-of-the-art VLMs on VMCBench, setting a new standard for scalable, consistent, and reproducible VLM evaluation.
@@ 2 Related Works
Figure 2: Challenges in evaluating open-ended questions. (Left) Rule-based metrics significantly underestimate model performance and penalize models that do not strictly follow the expected format. (Right) Model-based evaluations using two different versions of GPT yield substantially different scores, making comparisons inconsistent and raising reproducibility issues.
@@ 4.4 AutoConverter Results
Each agent in AutoConverter contributes to jointly improving correctness and difficulty. To analyze each agent’s role in AutoConverter, we conduct an ablation study by removing each component, as shown in Table 3(b). First, removing specialized error proposers results in a decrease in difficulty, with an average increase of 1.6% in relative accuracy. Next, omitting the reviewer for iterative distractor refinement leads to a 4.9% relative increase in average VLM performance, indicating lower difficulty. Finally, removing the question evaluator and refiner, which serve as correctness safeguards, causes an 8.7% relative decrease in correctness score. We also compare AutoConverter with a naive baseline that uses the prompt “Please generate 3 distractors given the question, answer, and image.” As shown in Table 3(b), AutoConverter significantly outperforms this baseline in both correctness (2.2% relative increase) and difficulty (11.5% relative decrease in average VLM performance), demonstrating the effectiveness of our approach.
@@ 5.1 Benchmark Overview
To ensure the correctness of the converted questions, AutoConverter includes an internal correctness evaluator that assigns a correctness score to each converted question, ensuring only one correct answer exists. Out of the 9,450 converted questions, 103, 162, 399, 635, and 8,151 questions received correctness scores of 1, 2, 3, 4, and 5, respectively. Six PhD student annotators reviewed all questions with correctness scores below 5 and randomly sampled 1,101 questions with a score of 5, resulting in a total of 2,400 annotated questions. The correctness rates for questions with scores of 1, 2, 3, 4, and 5 were 51%, 51%, 63%, 84%, and 95%, respectively. This high accuracy of the correctness evaluation supports the reliability of questions with a confidence score of 5 for future dataset construction. Additionally, we annotated the source of errors, distinguishing between errors from the ground-truth answers (original dataset errors) and those introduced by AutoConverter. Notably, 59% of errors were due to issues in the ground-truth answers, indicating that if all original open-ended questions were correct, AutoConverter would introduce only 2% errors in total for questions with a correctness score of 5. We removed all questions that were identified as incorrect, resulting in a final total of 9,018 questions.
@@ 5.2 Evaluation Results
We evaluated 33 state-of-the-art vision language models (VLMs) on the VMCBench test set to assess their performance, including GPT-4 [40], Gemini-1.5 [51], Claude-3.5 [2], Qwen2-VL [54], Molmo [8], Cambrian [52], VILA [26], CogVLM [16], Phi [1], Idefics2 [22], DeepSeek-VL [29], PaliGemma [3], LLaVA1.5 [27], Chameleon [50], and InstructBLIP [7]. The evaluation is conducted in the zero-shot setting with prompt “Question: {question} Options: A. {A} B. {B} C. {C} D. {D} Answer with the option’s letter from the given choices directly.” The results are summarized in Table 1. Detailed results are in Appendix.
@@
Private models remain the top performers, but the gap is narrowing. As shown in Table 1, the best-performing model on VMCBench is GPT-4o, achieving 80.6% overall accuracy. However, the gap between GPT-4o and the second-best model, Qwen2-VL-7B, is only 2.0%. This open-source model demonstrates that the performance difference between private and public models is decreasing.
Rapid advances in VLM development. Another notable trend is the significant improvement from InstructBLIP to Qwen2-VL-7B, nearly doubling in performance (40.9% vs. 78.6%). These models were released in 2023 and 2024, respectively, illustrating the rapid pace of progress in VLMs.
Scaling up models generally improves performance. As expected, scaling up model sizes typically leads to performance gains. For instance, across different model families (such as VILA, Cambrian, Chameleon, DeepSeek, and LLaVA), we observe a clear improvement with larger models, similar to scaling trends in language models [18, 4].
Evaluating across diverse datasets is essential to uncovering model limitations. A comparison between Phi-3 and Phi-3.5 reveals that while Phi-3.5 achieves better performance on some reasoning datasets such as MMMU and MathVista, it performs significantly worse on OCR and document/chart understanding tasks, which reduces its overall performance. This underscores the importance of a holistic evaluation of models across a broad range of datasets. Previously, this was challenging due to variations in dataset formats and evaluation protocols. VMCBench addresses this by unifying the task into a multiple-choice format, thus reducing evaluation costs and standardizing comparisons.
Continuous feature inputs outperform discrete tokens. Chameleon is the only model in our experiments that utilizes VQGAN discrete tokens, yet it performs significantly worse than smaller models trained with continuous features using less computational resources. This observation raises an intriguing question: why might VQGAN tokens be less effective for image understanding tasks?
@@ 6 Conclusion
We address limitations in open-ended visual question answering benchmarks for evaluating vision language models (VLMs) by introducing AutoConverter, a multi-agent system that transforms questions into multiple-choice format with challenging distractors. AutoConverter effectively creates questions that match or exceed the difficulty of human-written distractors, with VLMs often achieving similar or lower accuracy. Building on this, VMCBench provides a benchmark of 9,018 multiple-choice questions, setting a new standard for consistent and rigorous VLM evaluation.
@@ C.3 Scaling Trends on VMCBench
C.4 Additional Examples from VMCBench
To provide a deeper understanding of VMCBench, we include 60 examples (three from each of the 20 sources) in Tables 5 through 11. These examples demonstrate the high quality of our dataset and offer valuable insights for readers.

### v2

@@ Abstract
The rapid development of vision language models (VLMs) demands rigorous and reliable evaluation. However, current visual question answering (VQA) benchmarks often depend on open-ended questions, making accurate evaluation difficult due to the variability in natural language responses. To address this, we introduce AutoConverter, an agentic framework that automatically converts these open-ended questions into multiple-choice format, enabling objective evaluation while reducing the costly multiple-choice question creation process. Our experiments demonstrate that AutoConverter can generate correct and challenging multiple-choice questions, with VLMs demonstrating consistently similar or lower accuracy on these questions compared to human-created ones. Using AutoConverter, we construct VMCBench, a benchmark created by transforming 20 existing VQA datasets into a unified multiple-choice format, totaling 9,018 questions. We comprehensively evaluate 33 state-of-the-art VLMs on VMCBench, setting a new standard for scalable, consistent, and reproducible VLM evaluation.
@@ 2 Related Works
Figure 2: Challenges in evaluating open-ended questions. (Left) Rule-based metrics significantly underestimate model performance and penalize models that do not strictly follow the expected format. (Right) Model-based evaluations using two different versions of GPT yield substantially different scores, making comparisons inconsistent and raising reproducibility issues. The repeated points represent different model sizes within the same model family (e.g., GPT-4o/GPT-4o-Mini for 2 points).
@@ 4.4 AutoConverter Results
Each agent in AutoConverter contributes to jointly improving correctness and difficulty. To analyze each agent’s role in AutoConverter, we conduct an ablation study by removing each component, as shown in Figure 3(b). First, removing specialized error proposers results in a decrease in difficulty, with an average increase of 1.6% in relative accuracy. Next, omitting the reviewer for iterative distractor refinement leads to a 4.9% relative increase in average VLM performance, indicating lower difficulty. Finally, removing the question evaluator and refiner, which serve as correctness safeguards, causes an 8.7% relative decrease in correctness score. We also compare AutoConverter with a naive baseline that uses the prompt “Please generate 3 distractors given the question, answer, and image.” As shown in Figure 3(b), AutoConverter significantly outperforms this baseline in both correctness (2.2% relative increase) and difficulty (11.5% relative decrease in average VLM performance), demonstrating the effectiveness of our approach.
@@ 5.1 Benchmark Overview
To ensure the correctness of the converted questions, AutoConverter incorporates an internal correctness evaluator that assigns a correctness score to each converted question, ensuring that only one correct answer exists. Among the 9,450 converted questions, 103, 162, 399, 635, and 8,151 questions received correctness scores of 1–5, respectively. Six PhD student annotators reviewed all questions with correctness scores below 5 and randomly sampled 1,101 questions with a score of 5, resulting in a total of 2,400 annotated questions. The correctness rates for questions with scores of 1–5 were 51%, 51%, 63%, 84%, and 95%, respectively. This increasing trend in the correctness rate demonstrates the reliability of the correctness evaluation.
To gain further insight, we annotated the source of errors, distinguishing between those stemming from ground-truth answer inaccuracies (errors in the original dataset) and those introduced by AutoConverter. Our analysis revealed that 59% of errors were due to issues in the ground-truth answers rather than AutoConverter. Notably, if all original open-ended questions were error-free, AutoConverter would have introduced only 2% errors among questions with a correctness score of 5, highlighting the reliability of our approach. Therefore, we recommend retaining only questions with a correctness score of 5 for future benchmark construction. We removed all questions identified as incorrect, resulting in a final total of 9,018 questions.
@@ 5.2 Evaluation Results
We evaluated 33 state-of-the-art VLMs on the VMCBench, including GPT-4 [40], Gemini-1.5 [51], Claude-3.5 [2], Qwen2-VL [54], Molmo [8], Cambrian [52], VILA [26], CogVLM [16], Phi [1], Idefics2 [22], DeepSeek-VL [29], PaliGemma [3], LLaVA1.5 [27], Chameleon [50], and InstructBLIP [7]. The evaluation is conducted in the zero-shot setting with the prompt “Question: {question} Options: A. {A} B. {B} C. {C} D. {D} Answer with the option’s letter from the given choices directly.” The results are summarized in Table 1. Detailed results are in Appendix.
@@
The gap between private and public models is narrowing. As shown in Table 1, the best-performing model on VMCBench is the publicly available Qwen2-VL-72B, which achieves 85.0% overall accuracy—surpassing the best private model, GPT-4o, which scores 80.3%.
VLM development is advancing rapidly. Another notable trend is the substantial performance gain from InstructBLIP-7B to Qwen2-VL-72B, with accuracy nearly doubling from 42.1% to 85.0%. These models were released in 2023 and 2024, respectively, highlighting the rapid pace of progress in VLMs.
Scaling up models improves performance. As expected, scaling up model sizes leads to performance gains. For instance, across different model families (such as Qwen, Molmo, VILA, Cambrian, DeepSeek, and LLaVA), we observe a clear improvement with larger models, similar to scaling trends in language models [18, 4].
Diverse datasets reveal model limitations. While Phi-3.5 outperforms Phi-3 on reasoning tasks like MMMU and MathVista, it underperforms on OCR and doc/chart tasks, lowering its overall accuracy. This highlights the need for broad evaluation. VMCBench enables this by unifying diverse datasets into a standardized multiple-choice format.
@@ 6 Conclusion
We address the limitations of open-ended visual question answering benchmarks for evaluating vision language models (VLMs) by introducing AutoConverter, a multi-agent system that transforms questions into multiple-choice format with challenging distractors. AutoConverter generates questions that rival or surpass human-written ones in difficulty. Using it, we construct VMCBench, a benchmark of 9,018 multiple-choice questions that sets a new standard for consistent and rigorous VLM evaluation.
@@ Acknowledgements
Broader Impacts and Ethics Statement
AutoConverter and VMCBench simplify and standardize vision language model (VLM) evaluation, providing a valuable tool for advancing VLM development through scalable, consistent, and reproducible evaluation. Beyond VLM evaluation, AutoConverter can be applied to other domains, such as education, to generate challenging and high-quality questions. However, human verification is essential to ensure the generated questions align with their intended purpose, maintain proper value, and achieve the appropriate level of difficulty without introducing biases.
Reproducibility Statement
We provide an open-source implementation of our work at https://github.com/yuhui-zh15/AutoConverter. This will enable researchers to reproduce all the experiments in paper and conduct their own analyses.
Limitations
AutoConverter achieves high question correctness, but 5% of the questions with the highest correctness scores still contain errors, although half of which are due to flaws in the original datasets. Moreover, while VMCBench includes a diverse range of models and datasets, its coverage remains incomplete. Moving forward, we aim to leverage AutoConverter to expand dataset and model coverage, addressing evolving needs in VLM evaluation.
Summary of Appendix
  • In §A, we present further analysis of the evaluation failures for open-ended questions discussed in §3.
  • In §B, we provide supplementary details on the AutoConverter framework described in §4.
  • In §C, we elaborate more details of VMCBench as outlined in §5.
@@ C.3 Scaling Trends on VMCBench
C.4 Option Permutation Sensitivity on VMCBench
To assess whether VLMs are sensitive to option permutation, we randomly shuffled all correct options twice using two different random seeds. We observed no significant difference in performance on VMCBench (Table 2). This result demonstrates the robustness of multiple-choice evaluation in VMCBench.
Question $M_{1}$ $M_{2}$ $M_{3}$ $M_{4}$ $M_{5}$ $M_{6}$
Original 59.7 52.6 68.3 78.8 78.1 80.8
Shuffled 1 59.8 52.7 68.0 78.6 77.7 81.0
Shuffled 2 59.9 52.5 67.9 78.7 78.2 80.3
Table 2: Performance on VMCBench with shuffled options. $M_{1}$ is PaliGemma-3B, $M_{2}$ is LLaVA-1.5-7B, $M_{3}$ is Phi-3.5-Vision, $M_{4}$ is Qwen2-VL-7B, $M_{5}$ is Claude-3.5-Sonnet, $M_{6}$ is GPT-4o.
C.5 Additional Examples from VMCBench
To provide a deeper understanding of VMCBench, we include 60 examples (three from each of the 20 sources) in Tables 6 through 12. These examples demonstrate the high quality of our dataset and offer valuable insights for readers.